# Cosmological gravity on all scales IV: 3x2pt Fisher forecasts for pixelised phenomenological modified gravity

### v1

@@ 2.2 Computing the non-linear matter power spectrum with ReACT
The principle behind the halo model reaction formalism [42, 30, 43] is that one can greatly improve the accuracy in the halo model in the transition regime between the 2-halo term and the 1-halo term by computing the ratio dubbed the ‘non-linear response of the power spectrum’ or ‘reaction’ [44] i.e., the ratio of matter power spectra between two cosmologies that share the same linear power spectrum. The halo model reaction formalism computes the non-linear reaction for beyond $\Lambda$CDM cosmologies, given by the ratio of $P(k)$ in the target cosmology relative to the ‘pseudo spectrum’, which is the $\Lambda$CDM power spectrum with the the initial conditions modified to match the linear power spectrum in the modified gravity model at a specific target redshift. The code that implements this formalism, ReACT has been applied to phenomenological extensions of $\Lambda$CDM with specific predictions for the functional form of $\mu(z,k)$ [45, 46]. In the case where theoretical predictions for $\mu$ are derived from the effective field theory of dark energy, the non-linear recipe in ReACT was validated against emulators for $w$CDM, $f(R)$, $n$DGP cosmologies to within sub-percent accuracy.
@@ 2.3 Scale cuts with the BNT transformation
We will present our results for three different choices for $k_{\rm max}$ that capture the extent to which one trusts the modelling of the non-linear scales in the analysis. For the un-transformed $C_{ij}(\ell)$, we will use the same criteria for the galaxy clustering and galaxy-galaxy-lensing data, but choose appropriate values of $\ell_{\rm max}$ for the shear data. We compute the convariance matrix for the $k$-cut statistics included in the $3\times 2$pt data vector via the likelihood sampling method described in section III E of [49].
@@ 3.1 Impact of BNT transform: mitigating baryonic effects
One of the biggest challenges facing large scale structure in the coming decades is baryonic feedback. In order to quantify the effect of baryons on parameter constraints, we compute the Fisher matrix for two cases: one where we fix the AGN feedback parameter $T_{\rm AGN}$ to its fiducial value and one where we marginalize over it. We then repeat this procedure including the BNT transform in the modelling. By comparing the parameter constraints for each case (no BNT and for the BNT transformed data vector) with a scale cut of $k_{\rm cut}=0.5\,h\,{\rm Mpc}^{-1}$. We present the results of this analysis in fig. 4.
@@
We find that when we repeat this procedure for $k_{\rm cut}=1\,h\,{\rm Mpc}^{-1}$, the error on the parameters increases to up to 100%. These results indicate that phenomenological modified gravity is not sensitive to baryonic feedback when the BNT transform is employed for cosmic shear assuming $k_{\rm cut}-0.5\,h\,{\rm Mpc}^{-1}$.
@@ 3.2 Binning Schemes
Our initial study divides the redshift range in five bins in the range $0\leq z\leq 3$ (see table  1). Without loss of generality, we restrict ourselves here to the BNT transformed data vector with $k_{\rm cut}=0.5\,h\,{\rm Mpc}^{-1}$ (we check and find that these results are robust to the choice of $k_{\rm cut}$ and the choices of including BNT/concentration fit in the modelling). We find that for these choices, we are able to constrain almost all modified gravity parameters to within 10% (we will return to this point and quantitatively describe and discuss our results in the coming subsections), except for $\mu_{5}$ and $\eta_{5}$ ($2.15\leq z\leq 3$). The error on this bin is degraded by over an order of magnitude relative to the other bins. This is due to the fact that the structure formation is linear in this redshift range on the scales that enter our Fisher forecast. As we are focussing on the impact of non-linear structure formation, we therefore fix $\mu=\eta=1$ in this bin for the rest of this work.
@@
We now turn towards identifying the optimal number of bins for an LSST Y10-like survey. For this, we divide the redshift range between $0\leq z\leq 2.15$ into $N$ bins where $N=\{3,4,5,6,7\}$. The bin-widths correspond to equal $D(z)$ (as computed in $\Lambda$CDM). Since the concentration fitting function has been validated for $0.8\leq\mu\leq 1.2$ and $1.26\leq D(z)\leq 1.45$, it cannot be included in the modelling for $N\geq 5$ bins since the corresponding $D(z)<1.26$. We use the Fisher FOM and the average constraining power on the MG parameters to statistically compare different binning schemes. We show the results of this procedure on fig. 6. We find that we obtain the best modified gravity constraints for $N=4$ case, but the Fisher FOM for the $N=3$ bin case is larger due to the fact that the $\Lambda$CDM parameters are constrained better for this scheme. Since we find relatively good constraints on the cosmological parameters (see sec. 3.3 for more details), we will adopt the $N=4$ case as our ‘best’ case and will present our final results for this choice. We leave a detailed analysis of whether there is an even more optimal choice for the bin sizes and spacing to future work.
@@ 3.3 Impact of modelling on constraining power
Figure 8: We present the $1-\sigma$ and $2-\sigma$ contours on the cosmological parameters and the modified gravity parameters for the three different modelling cases described in the text (see table 3). We see that including the additional information from the concentration-mass fit and the greater control of the scales that we include in the modelling allows us to not only have greater constraining power, but also breaks degeneracies between parameter combinations.
@@ 3.4 Influence of degeneracies on modelling choices and binning strategies
As we mentioned before, irrespective of the BNT/concentration modelling, the most prominent degeneracy in fig. 8 occurs between the MG parameters in the same redshift bin $\mu_{i}$ and $\eta_{i}$ (along the diagonal). This is because the lensing observable is shifted up identically on linear scales when $\eta-1=2(\mu-1)$, which can be reconciled to the weighting factor $\Sigma=\mu(1+\eta)$ in the lensing $C(\ell)$ (see eq. (2.7)). This degeneracy therefore is stronger at low redshift, since the lensing signal dominates the $3\times 2pt$ data vector at low redshift (the clustering kernel is effectively zero at $z<0.2$). Of course, as we go from low to high redshift, the improved constraining power around $z\sim 1$ (since this where the kernels for both lensing and clustering peak) and the combined power $3\times 2$pt data is more effective. In particular, when modified gravity effects have sufficiently impacted structure formation by that time we have in particular.
@@
In addition to this diagonal degeneracy, we also note that the only off-diagonal degeneracy that persists for all our modelling cases is $\mu_{3}-\mu_{4}$. In general, applying the concentration on top of BNT only affects the degeneracies associated with the first two modified gravity bins. For our preferred case of BNT+conc+$k_{\rm{cut}}=0.5h$Mpc-1, there are small degeneracies between the modified gravity parameters and the LCDM parameters, which reduce the constraints on the LCDM parameters. The largest effect is for $\Omega_{m}$ for which the 1-$\sigma$ error increases from about 0.68% without modified gravity to 1% with modified gravity. We note that the LCDM parameters are all strongly constrained even when modified gravity is included. We also note that another advantage of the non-linear modelling and thus using a scale cut of $0.5h$Mpc-1, is that the degeneracies with LCDM are much smaller than in linear theory.
We now turn towards the choice of parametrisation. If one were to re-parameterise from $(\mu,\eta)\rightarrow(\mu,\Sigma)$, the degeneracy in the lensing $Cls$ may be seen explicitly in the Fisher contours where $\mu_{1},\mu_{2}$ and $\Sigma_{1},\Sigma_{2}$ are almost perfectly degenerate, respectively, on linear scales. In this case, the constraints on $\mu_{i}$ and $\Sigma_{i}$ are comparable, as opposed to $\mu-\eta$ where $\mu$ is generally constrained better than $\eta$ (see appendix A for more detail). Since the basis of the non-linear recipe in our work is in the validation of $P(k)$ with $N$-body simulations for $0.8\leq\mu\leq 1.2$, and the resultant constraints on $\mu$ are stronger, we choose the $\mu-\eta$ parameterisation to present all of our results. This also ensures the future combined analyses with spectroscopic surveys yield optimal results. We discuss in appendix A details on the Fisher constraints for the $\mu-\Sigma$ parameterisation.
@@ A.2 Fisher contours
We now turn to the subject of alternative parameterisations. We show in fig. 12 the Fisher contours associated to the $\mu-\Sigma$ parameterisation. We explicitly see here the $\mu-\Sigma$ degeneracy discussed in sec. 3.4 and the overall change in the constraining power on the MG parameters relative to the standard $\mu-\eta$ case considered in the main text. In the $\mu-\Sigma$ case, the transformation homogenises the constraints in the MG parameter space, i.e., the relative error on $\mu$ is comparable to that of $\Sigma$, as opposed to $\mu$ and $\eta$ (see sec. 3.4 for more details). In the main text, we made the choice of presenting our results using the $\mu-\eta$ parameterisation, since it leads to better constraints on $\mu$, which is the parameter that requires our non-linear recipe and this is varied in the $N$-body simulations used to validate our non-linear recipe. We note that the specific probe combination being considered is particularly relevant for this choice. Galaxy clustering constrains $\mu$ alone, while weak-lensing directly constrains $\Sigma=\mu(1+\eta)$. The nature of the data in question (spectroscopic/photometric) is also relevant to the choice of parameterisation, particularly becayse spectroscopic surveys will not be able to effectively constrain $\eta/\Sigma$. We conclude that these considerations make it difficult to make an overall blanket statement about which choice is better in general. Instead, we suggest that these results indicate what one can expect for a typical photometric survey, and may be used as a guide to make an informed decision depending on the goals of the analysis one is interested in.

### v2

@@ 2.2 Computing the non-linear matter power spectrum with ReACT
The principle behind the halo model reaction formalism [ref:ReactTheory, 42, 30, 43] is that one can greatly improve the accuracy in the halo model in the transition regime between the 2-halo term and the 1-halo term by computing the ratio dubbed the ‘non-linear response of the power spectrum’ or ‘reaction’ [44] i.e., the ratio of matter power spectra between two cosmologies that share the same linear power spectrum. The halo model reaction formalism computes the non-linear reaction for beyond $\Lambda$CDM cosmologies, given by the ratio of $P(k)$ in the target cosmology relative to the ‘pseudo spectrum’, which is the $\Lambda$CDM power spectrum with the the initial conditions modified to match the linear power spectrum in the modified gravity model at a specific target redshift. The code that implements this formalism, ReACT has been applied to phenomenological extensions of $\Lambda$CDM with specific predictions for the functional form of $\mu(z,k)$ [45, 46]. In the case where theoretical predictions for $\mu$ are derived from the effective field theory of dark energy, the non-linear recipe in ReACT was validated against emulators for $w$CDM, $f(R)$, $n$DGP cosmologies to within sub-percent accuracy.
@@ 2.3 Scale cuts with the BNT transformation
We will present our results for three different choices for $k_{\rm max}$ that capture the extent to which one trusts the modelling of the non-linear scales in the analysis. For the un-transformed $C_{ij}(\ell)$, we will use the same criteria for the galaxy clustering and galaxy-galaxy-lensing data, but choose appropriate values of $\ell_{\rm max}$ for the shear data. We compute the covariance matrix for the $k$-cut statistics included in the $3\times 2$pt data vector via the likelihood sampling method described in section III E of [49].
@@ 3.1 Impact of BNT transform: mitigating baryonic effects
One of the biggest challenges facing large scale structure in the coming decades is baryonic feedback. In order to quantify the effect of baryons on parameter constraints, we compute the Fisher matrix for two cases: one where we fix the AGN feedback parameter $T_{\rm AGN}$ to its fiducial value and one where we marginalize over it. We then repeat this procedure including the BNT transform in the modelling. By comparing the parameter constraints with and without marginalisation, and with and without the BNT transformation, we can see the effect that uncertainty in the baryon modelling has on our analysis in each case. We set $k_{\rm cut}=0.5\,h\,{\rm Mpc}^{-1}$ initially, and comment on relaxing this below. We present the results of this analysis in fig. 4.
@@
We find that when we repeat this procedure for $k_{\rm cut}=1\,h\,{\rm Mpc}^{-1}$, the error on the parameters increases to up to 100%. In addition, the constraints in table 3 show that no parameters have worse constraints when the BNT transform is used. Combined, these results indicate that the BNT transform with $k_{\rm cut}-0.5\,h\,{\rm Mpc}^{-1}$ should be used when doing a phenomenological modified gravity analysis with stage IV data, and that when doing so the analysis is not sensitive to baryonic feedback.
@@ 3.2 Binning Schemes
Our initial study divides the redshift range in five bins in the range $0\leq z\leq 3$ (see table  1). Without loss of generality, we restrict ourselves here to the BNT transformed data vector with $k_{\rm cut}=0.5\,h\,{\rm Mpc}^{-1}$ (we check and find that these results are robust to the choice of $k_{\rm cut}$ and the choices of including BNT/concentration fit in the modelling). We find that for these choices, we are able to constrain almost all modified gravity parameters to within 10% (we will return to this point and quantitatively describe and discuss our results in the coming subsections), except for $\mu_{5}$ and $\eta_{5}$ ($2.15\leq z\leq 3$). The error on this bin is degraded by over an order of magnitude relative to the other bins. This is due to the fact that the structure formation is linear in this redshift range on the scales that enter our Fisher forecast. As we are focusing on the impact of non-linear structure formation, we therefore fix $\mu=\eta=1$ in this bin for the rest of this work.
@@
We now turn towards identifying the optimal number of bins for an LSST Y10-like survey. For this, we divide the redshift range between $0\leq z\leq 2.15$ into $N$ bins where $N=\{3,4,5,6,7\}$. The bin-widths correspond to equal $D(z)$ (as computed in $\Lambda$CDM). Since the concentration fitting function has been validated for $0.8\leq\mu\leq 1.2$ and $1.26\leq D(z)\leq 1.45$, it cannot be included in the modelling for $N\geq 5$ bins since the corresponding $D(z)<1.26$. We use the Fisher FOM and the average constraining power on the MG parameters to statistically compare different binning schemes. We show the results of this procedure on fig. 6. We find that we obtain the best modified gravity constraints for $N=4$ case, but the Fisher FOM for the $N=3$ bin case is larger due to the fact that the $\Lambda$CDM parameters are constrained better for this scheme. Since we find relatively good constraints on the cosmological parameters (see sec. 3.3 for more details), we will adopt the $N=4$ case as our ‘best’ case and will present our final results for this choice. [^3] Note that we repeated this analysis for the 3 and 4 bin cases whilst including the effect of the concentration, and the qualitative results did not change. We leave a detailed analysis of whether there is an even more optimal choice for the bin sizes and spacing to future work.
@@ 3.3 Impact of modelling on constraining power
Figure 8: We present the $1-\sigma$ and $2-\sigma$ contours on the cosmological parameters and the modified gravity parameters for the middle and best different modelling cases described in the text (see table 3). We see that including the additional information from the concentration-mass fit and the greater control of the scales that we include in the modelling allows us to not only have greater constraining power, but also breaks degeneracies between parameter combinations.
@@ 3.4 Influence of degeneracies on modelling choices and binning strategies
As we mentioned before, irrespective of the BNT/concentration modelling, the most prominent degeneracy in fig. 8 occurs between the MG parameters in the same redshift bin $\mu_{i}$ and $\eta_{i}$ (along the diagonal). This is because the lensing observable is shifted up identically on linear scales when $\eta-1=2(\mu-1)$, which can be reconciled to the weighting factor $\Sigma=\mu(1+\eta)$ in the lensing $C(\ell)$ (see eq. (2.7)). This degeneracy therefore is stronger at low redshift, since the lensing signal dominates the $3\times 2pt$ data vector at low redshift (the clustering kernel is effectively zero at $z<0.2$). As we go from low to high redshift, although the lensing itself increases in constraining power (since this is where the kernels for both lensing and clustering peak), the relative importance of the clustering increases as well, resulting in the degeneracy being weaker. This can be seen in fig. 8 where the 2D $\mu_{i}$-$\eta_{i}$ ellipse is narrower and more elongated for $i=1,2$ than for $i=3,4$.
@@
In addition to this diagonal degeneracy, we also note that the only off-diagonal degeneracy that persists for all our modelling cases is $\mu_{3}-\mu_{4}$. In general, as can be seen in fig. 8, applying the concentration on top of BNT only affects the degeneracies associated with either of the first two modified gravity bins ($\mu_{1-2}$ and $\eta_{1-2}$). For our preferred case of BNT+concentration+$k_{\rm{cut}}=0.5\,h\,$Mpc-1, there are small degeneracies between the modified gravity parameters and the $\Lambda$CDM parameters, which reduce the constraints on the $\Lambda$CDM parameters. The largest effect is for $\Omega_{m}$ for which the 1-$\sigma$ error increases from about 0.68% without modified gravity to 1% with modified gravity. We note that the $\Lambda$CDM parameters are all strongly constrained even when modified gravity is included. We also note that another advantage of the non-linear modelling and thus using a scale cut of $0.5h$Mpc-1, is that the degeneracies with $\Lambda$CDM are much smaller than in linear theory.
We now turn towards the choice of parameterisation. If one were to re-parameterise from $(\mu,\eta)\rightarrow(\mu,\Sigma)$, the degeneracy in the lensing $Cls$ may be seen explicitly in the Fisher contours where $\mu_{1},\mu_{2}$ and $\Sigma_{1},\Sigma_{2}$ are almost perfectly degenerate, respectively, on linear scales. In this case, the constraints on $\mu_{i}$ and $\Sigma_{i}$ are comparable, as opposed to $\mu-\eta$ where $\mu$ is generally constrained better than $\eta$ (see appendix A for more detail). Since the basis of the non-linear recipe in our work is in the validation of $P(k)$ with $N$-body simulations for $0.8\leq\mu\leq 1.2$, and the resultant constraints on $\mu$ are stronger, we choose the $\mu-\eta$ parameterisation to present all of our results. This also ensures the future combined analyses with spectroscopic surveys yield optimal results. We discuss in appendix A details on the Fisher constraints for the $\mu-\Sigma$ parameterisation.
@@ A.2 Fisher contours
We now turn to the subject of alternative parameterisations. We show in fig. 12 the Fisher contours associated to the $\mu-\Sigma$ parameterisation. We explicitly see here the $\mu-\Sigma$ degeneracy discussed in sec. 3.4 and the overall change in the constraining power on the MG parameters relative to the standard $\mu-\eta$ case considered in the main text. In the $\mu-\Sigma$ case, the transformation homogenises the constraints in the MG parameter space, i.e., the relative error on $\mu$ is comparable to that of $\Sigma$, as opposed to $\mu$ and $\eta$ (see sec. 3.4 for more details). In the main text, we made the choice of presenting our results using the $\mu-\eta$ parameterisation, since it leads to better constraints on $\mu$, which is the parameter that requires our non-linear recipe and this is varied in the $N$-body simulations used to validate our non-linear recipe. We note that the specific probe combination being considered is particularly relevant for this choice. Galaxy clustering constrains $\mu$ alone, while weak-lensing directly constrains $\Sigma=\mu(1+\eta)$. The nature of the data in question (spectroscopic/photometric) is also relevant to the choice of parameterisation, particularly because spectroscopic surveys will not be able to effectively constrain $\eta/\Sigma$. We conclude that these considerations make it difficult to make an overall blanket statement about which choice is better in general. Instead, we suggest that these results indicate what one can expect for a typical photometric survey, and may be used as a guide to make an informed decision depending on the goals of the analysis one is interested in.